A nurse must add a 5% solution to 750 ml of a 30% to produce a 20% solution. How much of the 5% solution must be added?
Answer by abhinav75(8)

let x volume is added...
initial volume is 750 ml = 0.75 litres
5x+ 30(0.75) = 20(x+0.75)
=> 5x + 22.5 = 20x + 15
=> 15x = 7.5
=> x = 0.5
therefore, 500 ml should be added of the 5% solution